OpenStudy (anonymous):

Solve the system of equations. 2x – 9y = 32 4x + y = -12

OpenStudy (anonymous):

well...you can try to multiply either the x or y with a certain number so that atleast one can cancel...for instance, if you want to cancel the x multiply the 2x - 9y = 32 by -2

OpenStudy (anonymous):

can you go from there?

OpenStudy (anonymous):

x= -2 y=-4

OpenStudy (zepp):

Do you know how the reduction method works?

OpenStudy (anonymous):

2x-9y=32.......................1 4x+y=-12..................2 from equ 1 2x=32+9y x=32+9y/2 now put these values in 2nd equ then solve.

OpenStudy (zepp):

2x – 9y = 32 4x + y = -12 We have these two equations here, we want to make a variable disappear when added/substracted I chose to add them up, I don't want the y, so I multiply the second equation by 9 (4x + y = -12)*9 36x + 9y = -108 Now I can add them 2x – 9y = 32 + 36x + 9y = -108 ============ 36x + 2x = 32 + -108 38x = 32 - 108 Can you do it now?

OpenStudy (anonymous):

add 32?

OpenStudy (kropot72):

You need to solve the following equation: \[38x=32-108\] The first step is to reduce the right hand side of the equation to a single number by adding 32 and -108 (the result will be a negative number). Can you do that?

OpenStudy (anonymous):

38x=140?

OpenStudy (kropot72):

As I said the result of adding 32 and -108 will be a negative number. We are adding a positive number and a negative number: 32 + (-108)=? Hint: 40 + (-100) = -60

OpenStudy (kropot72):

To find 32 - 108 you can subtract 32 from 108 and give the result a negative sign.

OpenStudy (anonymous):

2x-9y=32.......................1 4x+y=-12..................2 multiply 1 eua with 2 on both sides 4x-18y=64------------3 now equa2-eqau3 4x+y-(4x-18y)=-12-64 19y=-76 y=-4 use this y val inn eqa 1 we get 2x+36=32 2x=-2 x=-2